OpenStudy (anonymous):

find the equation for the line (-4,6) and parallel to y=-3x+4

OpenStudy (anonymous):

Well the sloped have to be equal to be parallel. (y-y1)=m(x-x1) (y-6)=-3(x--4) y=-3x-12+6 y=-3x-6 -3x-y-6=0

OpenStudy (anonymous):

thanks boo

OpenStudy (anonymous):

No problem.